Yahaira:

 If you add​ Natalie's age and​ Fred's age, the result is 36. If you add​ Fred's age to 3 times​ Natalie's age, the result is 60. Write and solve a system of equations to find how old Fred and Natalie are.

kacchan:

Let Fred's age be y. Add​ Fred's age to 3 times​ Natalie's age, the result is.....

Constipated:

@XioGonz

mcSqueaks:

@supie

XioGonz:

x=Fred y=Natalie x+y=37 x+3y=63 2y+37=63 2y=63−37 2y=26 y=13 Now we know natalies age x+y=37 x+13=37 x=37−13 x=24 And that is Fred's age So there you go

Yahaira:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz x=Fred y=Natalie But they need to add to 36 not 37 x+y=37 x+3y=63 2y+37=63 2y=63−37 2y=26 y=13 Now we know natalies age x+y=37 x+13=37 x=37−13 x=24 And that is Fred's age So there you go \(\color{#0cbb34}{\text{End of Quote}}\)